two consecutive odd numbers total -20. Find the smaller.
Answer by checkley77(12569)

Let x & x=2 be the 2 numbers.
x+x+2=-20
2x=-20-2
2x=-22
x=-22/2
x=-11 ans. for the smaller number.
-11+2=-9 ans. for the other number.
Proof:
-11+(-9)=-20
-11-9=-20
-20=-20